OpenStudy (anonymous):

A particle passes through the point P = (4, 3, −5) at time t = 5, moving with constant velocity v = 4 i  − 3 j  + k. Find parametric equations for its motion. (Enter your answers as a comma-separated list of equations.)

OpenStudy (anonymous):

Hmph. What class is this for? I'm sure there's a prettier way to do this, but you can muscle through it by splitting it all up into its components... What have you got so far?

OpenStudy (anonymous):

\[ \textrm {Well, to get started, you know that your velocities in the} \\ \ \hat i , \hat j , \hat k \ \\ \textrm{ directions are all exclusive of each other. Hooray!} \\ \textrm{ So if you separate it into components, you get} \\ \ \\ v_x = 4 units/s \\ v_y = -3 units/s \\ v_z = 1 units/s \\ \ \\ \textrm{so with that you can set up three equations of motion;} \\ \textrm {we'll do it in one fell swoop to save time} \\ \ \\ \textbf v = \frac{d \textbf r }{dt} \\ \ \\ \int_0^{t} \textbf v dt = \int_{r_o}^r d \textbf r \\ \ \\ \textbf v t = \textbf r - \textbf r_o \\ \ \\ \textrm{which you can break up into the three equations for x, y, and z} \\ \ \\ v_x t = x - x_o \\ v_y t = y - y_o \\ v_z t = z - z_o \\ \ \\ \textrm{You can finally plug in all of the given values} \\ \textrm{to solve for your initial starting positions (xo, yo, zo), then} \\ \textrm{rearrange your eqs to be in the form } \\ \ \\ \\ \textbf r = \textbf v t + \textbf r_o \\ \ \\ \\ \textrm{There should be three of them; your solutions! } \]

OpenStudy (anonymous):

Oh my gosh, thank you so much! :) This is vector calculus by the way :p sorry I didn't reply sooner!

OpenStudy (anonymous):

Welcome, and none worries :)